OpenStudy (anonymous):

Urgent Help Please!! Of 450 college students, 120 are enrolled in math, 225 are enrolled in English, and 55 are enrolled in both. If a student is selected at random, find the probability of the following. (f) The student is not enrolled in English or is enrolled in mathematics.

OpenStudy (anonymous):

first you would add the students enrolled in each of the subjects, or both 225=120=55=400 450-400=50 the probability is 50/400, or 1/8, or in decimal form is .125

OpenStudy (kropot72):

The number of students not enrolled in English is 450 - 225 = 225. Therefore the probability that a randomly selected student is not enrolled in English is 225/450. The number of students enrolled in Math but not in English is 120 - 55 = 65. Therefore the probability that a randomly selected student is enrolled in Math but not in English is 65/450. the two events are mutually exclusive therefore the required probability is given by: \[\frac{225}{450}+\frac{65}{450}=you\ can\ calculate\]

OpenStudy (anonymous):

225+120+55=400*** by bad

OpenStudy (anonymous):

I calculated 290/450=.644 but i still keep getting the answer wrong am i doing something wrong?

OpenStudy (anonymous):

the question is worded weird, haha, is it trying to say the probability of somebody not enrolled in either subject/, bc that is what I solved for

OpenStudy (anonymous):

I believe it is asking not enrolled in English but enrolled in math

OpenStudy (anonymous):

oh okay, then I would say 120+55=175 175/450= appx. .389 percent of a chance

OpenStudy (anonymous):

Alright Thank You =)